What does y equal to in the equation 42y+5=17

Found 2 solutions by rfer, scampbell9972:
Answer by rfer(16322)

42y+5=17
42y=17-5
y=12/42
y=2/7

Answer by scampbell9972(1)

y = 0.285714286